OpenStudy (mtalhahassan2):

Find the exact non-decimal value of tan 27pi/6

OpenStudy (mtalhahassan2):

27pi/6 (divide by 3) ------- 9pi/2

OpenStudy (mtalhahassan2):

@Astrophysics

OpenStudy (mtalhahassan2):

@SolomonZelman @Shalante @Miracrown

OpenStudy (mtalhahassan2):

@ganeshie8 (Moderator)

OpenStudy (mtalhahassan2):

@ParthKohli

OpenStudy (astrophysics):

\[\frac{ 27 \pi }{ \left( \frac{ 6 }{ 3 } \right) }\]?

OpenStudy (mtalhahassan2):

no

OpenStudy (mtalhahassan2):

tan 27pi/6

OpenStudy (mtalhahassan2):

@ParthKohli the question is this

OpenStudy (astrophysics):

 here is the pi/6 triangle use it as a reference angle with the trig ratios

OpenStudy (mtalhahassan2):

what about 27?

OpenStudy (mtalhahassan2):

what happen to 27?

Parth (parthkohli):

\[\tan\left(4\pi + \frac{3\pi}{6}\right)\]\[= \tan\left(\frac{\pi}2\right)\]

OpenStudy (mtalhahassan2):

ok goted thnx